the ages of three brothers are x, 2x+1 and 3x-5. given that the sum of their ages is 38 find the age of the oldest brother
Answer by addingup(3677)

x+2x+1+3x-5 = 38
6x-4 = 38
6x = 42
x = 7
substitute this number for x and calculate:
x = 7
2x+1 = 2*7+1 =
3x-5 = 3*7-5 =
you should be able to do those calculations in your head.